a circuit with a 2 cell battery, and a resistor(R1) connected in series. Also 2 resistors in parallel at one point in the circuit(R2 & R3) and another 2 resistors in parallel at another point(R4 & R5). V1 = 4.0V, V2 = 1.5V , I2 = 0.2A , I1 = 0.7A, I4 = 0.3A, and R4 = 70Ξ©. Find all other voltages, currents, resistances and the totals for the circuit


1
Expert's answer
2021-04-08T09:02:38-0400

Explanations & Calculations


  • Refer to the attached figure.


  1. Since the current through the series resistor R1 represents the total current, then the total current is "\small I=i_1 =0.7A"
  2. Applying V=iR to R1,

"\qquad\qquad
\begin{aligned}
\small R_1&=\small \frac{V}{i_1}=\frac{4V}{0.7A}=\bold{5.7\Omega}
\end{aligned}"

3.Applying "\small V=iR" to R2,

"\qquad\qquad
\begin{aligned}
\small R_2&=\small \frac{V}{i_2}=\frac{1.5V}{0.2A}=\bold{7.5\Omega}
\end{aligned}"

4.Current flow through R3 is

"\qquad\qquad
\begin{aligned}
\small i_3&=\small I-i_2=0.7-0.2=\bold{0.5A}
\end{aligned}"

5.Since R2 & R3 are parallel to each other, they experience the same potential drop across them.

Therefore,

"\qquad\qquad
\begin{aligned}
\small V_3&=\small V_2=\bold{1.5V}
\end{aligned}"

6.Then applying "\small V=iR" to R3,

"\qquad\qquad
\begin{aligned}
\small R_3&=\small \frac{V}{i}=\frac{1.5V}{0.5A}=\bold{3\Omega}
\end{aligned}"

7.Applying "\small V=iR" to R4,

"\qquad\qquad
\begin{aligned}
\small V_4 &=\small iR=0.3A\times70\Omega=\bold{21V}
\end{aligned}"

8.Since R5 is connected in parallel to R4 they both experience the same potential drop. Then,

"\qquad\qquad
\begin{aligned}
\small V_5&=\small \bold{21V}
\end{aligned}"

9.The current through R5 is

"\qquad\qquad
\begin{aligned}
\small i_5&=\small 0.7-0.3=\bold{0.4A}
\end{aligned}"

10.Applying "\small V=iR" to R5,

"\qquad\qquad
\begin{aligned}
\small R_5&=\small \frac{V}{i}=\frac{21V}{0.4A}=\bold{52.5\Omega}
\end{aligned}"

11.Total voltage provided by the battery hence the potential of the battery.

"\qquad\qquad
\begin{aligned}
\small V_b&=\small 4V+1.5V+21V=\bold{26.5V}
\end{aligned}"

12.In case of an equivalent resistance that is external to the circuit,

"\qquad\qquad
\begin{aligned}
\small R_e&=\small R_1+\frac{R_2.R_3}{R_2+R_3}+\frac{R_4.R_5}{R_4+R_5}\\
&=\small 5.7+\frac{7.5\times3}{7.5+3}+\frac{70\times52.5}{70+52.5}\\
&=\small 5.7+2.14+30\\
&=\small \bold{37.84\Omega } 
\end{aligned}"

Which can also be calculated by

"\qquad\qquad
\begin{aligned}
\small R_e &=\small \frac{26.5V}{0.7A}=37.84\Omega
\end{aligned}"